OpenStudy (anonymous):

What are the coordinates of the vertex for f(x) = x2 + 4x + 10? (−2, 2) (2, 2) (−2, 6) (2, 6)

OpenStudy (anonymous):

Do you mean \[x^2+4x+10\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The answer would be: C, (-2,6) I just put it in my graphing calculator and it shows the vertex

OpenStudy (triciaal):

the x coordinate for the vertex is -b/2a given the eqn in the form ax^2 + bx + c the y value is f(x)

OpenStudy (triciaal):

-b/2a = -(4)/2*1 = -2 f(-2) = (-2)^2 + 4(-2) + 10 = 4 - 8 + 10 = 6 coordinates of the vertex using my natural calculator (-2, 6) as above